OpenStudy (anonymous):

What is the simplified form of the expression 4the square root of v. + 4 the square root of v to the sixth power. − the square root of the quantity 25 times v. − the square root of the quantity 9 times v to the sixth power.?

OpenStudy (anonymous):

i dont know where to start..

OpenStudy (anonymous):

\[4\sqrt{v}+4\sqrt{v^6}-\sqrt{25v}-\sqrt{9v^6}\]??

OpenStudy (anonymous):

like that? hard to read

OpenStudy (anonymous):

yes exactly that

OpenStudy (anonymous):

ok fine lets take it step by step

OpenStudy (anonymous):

\[\sqrt{v^6}\] is the square root of \(v^6\) since \(2\) goes in to 6 3 times, this is the same as \(v^3\)

OpenStudy (anonymous):

another way to see it is \[\sqrt{v^6}=\sqrt{(v^3)^2}=v^3\]

OpenStudy (anonymous):

now \(\sqrt{25v}=\sqrt{25}\times \sqrt{v}=5\sqrt{v}\)

OpenStudy (anonymous):

and finally since we already know \(\sqrt{v^6}=v^3\) we know \[\sqrt{9v^6}=3v^3\]

OpenStudy (anonymous):

now we have \[4\sqrt{v}+4v^3-5\sqrt{v}-3v^3\] and all that is left is to combine like terms

OpenStudy (anonymous):

\(4v^3-3v^3=v^3\) and \(4\sqrt{v}-5\sqrt{v}=-\sqrt{v}\) for a "final answer" of \[v^3-\sqrt{v}\]

OpenStudy (anonymous):

wow thats alot of steps, thank you so much, im going to go over and rad this a few times to fully unnderstand it, but hanks again so much!

OpenStudy (anonymous):

it is a lot of steps but that is largely because i was explaining as i went along, and skipped nothing with some practice you can do most of this in your head

OpenStudy (anonymous):

no steps were skipped, so if you really do read it, you will see all that you need to use to solve it

OpenStudy (anonymous):

oh, and your welcome!

OpenStudy (anonymous):

This man ( @satellite73 ) deserves a cookie :D lol